# Epistemic Diversity and Knowledge Collapse in Large Language Models

### v5

@@ Conclusions
Our results show that overall epistemic diversity is low in our pool of topics when compared to a baseline traditional search. That being said, for most model families and sizes tested, we observe an encouraging trend of knowledge expansion, indicating that so far, LLMs do not appear to be locked in to their narrow epistemic frames. Our results also highlight that RAG and the use of smaller models can help stave off knowledge collapse going forward. Going deeper, we find that the choice of RAG database may be important for improving epistemic representation across cultural contexts, as not all country-specific topics see equal gains. Here, practitioners need to be cautious about expanding their RAG sources with LLM-generated content. Finally, compared to a traditional knowledge source (Wikipedia) in both English and local languages for country-specific topics, we find that the claims generated in our study reflect English language knowledge more than local language knowledge, highlighting the need to investigate how local knowledge can be incorporated into LLM outputs. The general methodology presented in this paper can be used in the future to study epistemic diversity for any arbitrary set of topics, downstream tasks, and real-world use cases with open-ended plain-text LLM outputs. This allows researchers to answer research questions about which, whose, and how much knowledge LLMs are representing
@@ 5 Undisclosed August 2025 gpt-5-2025-08-07
A.4 Model Details
@@
A.5 InfoGap Details
We use GPT 5 to decompose English and local language Wikipedias to individual claims, retrieve candidate matches to LLM generated claims using a multilingual sentence embedding model, [^13] HuggingFace: sentence-transformers/LaBSE and predict if the retrieved multilingual claims match to the LLM generated English claims using GPT 5 Samir et al. (2024). We exclude claims generated using RAG to avoid Wikipedia text being included in the generation context.
A.6 Generation Details

### v6

@@ Conclusions
Our results show that overall epistemic diversity is low in our pool of topics when compared to a baseline traditional search, indicating a risk of knowledge collapse at the current time. On the other hand, for most model families and sizes tested we observe an encouraging trend of knowledge expansion, indicating that so far, LLMs do not appear to be locked in to narrow epistemic frames. Our results also highlight that RAG and the use of smaller models can help reduce the risk of knowledge collapse going forward. Going deeper, we find that the choice of RAG database may be important for improving epistemic representation across cultural contexts, as not all country-specific topics see equal gains. Here, practitioners need to be cautious about expanding their RAG sources with LLM-generated content, especially considering the negative impact to diversity of LLMs being recursively exposed to their own content Guo et al. (2024). Finally, compared to a traditional knowledge source (Wikipedia) in both English and local languages for country-specific topics, we find that the claims generated in our study reflect English language knowledge more than local language knowledge, highlighting the need to investigate how local knowledge can be incorporated into LLM outputs. The general methodology presented in this paper can be used in the future to study epistemic diversity for any arbitrary set of topics, downstream tasks, and real-world use cases with open-ended plain-text LLM outputs. This allows researchers to answer research questions about which, whose, and how much knowledge LLMs are representing
@@ 5 Undisclosed August 2025 gpt-5-2025-08-07
A.4 Comparison to Semantic Similarity Only
As described in § 3, an alternative approach to our methodology would be to measure diversity based solely on some notion of semantic similarity as performed in previous work Jiang et al. (2025); Wright et al. (2024). We now contrast our methodology with two recent popular approaches to diversity measurement which use semantic similarity, in order to demonstrate what is gained by focusing on semantic equivalence of claims as the unit of knowledge.
A.4.1 LLM Tropes Wright et al. (2024)
What about partitioning into classes based on semantic similarity? We now compare to the approach from Wright et al. (2024) which does exactly this. This is performed by taking the decomposed claims from our study, transforming them into sentence embeddings, and using DBSCAN Ester et al. (1996) to partition claims into unique classes. Note that DBSCAN parameters require tuning — for the sake of comparison, we selected reasonable parameters which resulted in a comparable number of clusters to the semantic equivalence case (eps of 0.2, min size of 2), but better results could potentially be achieved by hyperparameter tuning. After partitioning claims as such, we perform the same LLM-as-a-judge evaluation as in § 4.2, acquiring a cluster cohesion score of 3.48 (vs. 4.17 for our algorithm) and missing singleton rate of 0.4% (vs. 13.8% for our algorithm). As such, the recall of these clusters is higher (reflected in fewer singletons which should have been clustered), but the overall cohesion goes down, reflecting a tradeoff in terms of clustering quality.
A.4.2 Artificial Hivemind Jiang et al. (2025)
We now compare to the approach used in Jiang et al. (2025). In this work, diversity within LLM responses is measured using the average pairwise cosine similarity between responses represented as sentence embeddings. In other words, each LLM response is passed through a sentence embedding model (OpenAI’s text-embedding-3-small API in the original paper, Huggingface ID : all-MiniLM-L6-v2 here), and then the cosine similarity of each pair of embeddings is measured and averaged. We perform exactly this procedure on the original LLM responses of each model within each topic in our study, and plot diversity over time and across models (exactly as in Fig. 3) in Fig. 9.
[…]
A.5 Model Details
@@
A.6 InfoGap Details
We use GPT-5 to decompose English and local language Wikipedias to individual claims, retrieve candidate matches to LLM generated claims using a multilingual sentence embedding model, [^15] HuggingFace: sentence-transformers/LaBSE and predict if the retrieved multilingual claims match to the LLM generated English claims using GPT-5 Samir et al. (2024). We exclude claims generated using RAG to avoid Wikipedia text being included in the generation context.
A.7 Generation Details
@@
Appendix B Additional Figures and Tables
B.1 Empirical Comparison to Only Using Semantic Similarity
Figure 9: Average pairwise cosine similarity of within-topic model responses vs. model release date. Each point is a single model, with lines connecting models of approximately the same size across released versions. Error bars are 95% boostrapped confidence intervals based on cosine similarity within each topic (N=155). Lower scores mean greater diversity. While we see that search is still much more diverse than LLMs, many of the trends identified via epistemic diversity are lost.